# revtex4-2Repair the float

### v1

@@ Abstract
A new type of magnet called $p$-wave unconventional magnet is proposed, stimulated by the discovery of altermagnet. We study the tunneling conductance of $p$-wave unconventional magnet/superconductor junctions by adopting the effective Hamiltonian of $p$-wave unconventional magnets with time-reversal symmetry breaking, suggested in Ref [arXiv: 2309.01607 (2024)]. The tunneling conductance shows an asymmetric behavior with respect to bias voltage in the helical $p$-wave superconductor junctions. It is caused by the lack of helical edge states contributing to the charge conductance owing to the momentum-dependent spin-split feature of the Fermi surface in $p$-wave unconventional magnets. In chiral $d$ and $p$-wave superconductor junctions, the resulting spin-resolved tunneling conductance takes a different value for spin sectors due to the time-reversal symmetry breaking in superconductors. Our results qualitatively reproduce the results based on the simplified Hamiltonian in Ref [J. Phys. Soc. Jpn. 93, 114703 (2024)], where only the odd function of the exchange coupling of $p$-wave unconventional magnets is taken into account, which gives the shift of the Fermi surface and preserves the time-reversal symmetry similar to the spin-orbit coupling.
@@ I Introduction
Charge transport of superconducting junctions is a fundamental problem of superconductivity and mesoscopic physics due to the presence of Andreev reflection [1, 2]. Andreev reflection is a unique process in inhomogeneous superconductors (SCs), like superconducting junctions, where an injected electron (hole) is reflected as a hole (electron) at the interface, and it can generate the inner gap states called Andreev bound states [3, 4, 5]. In unconventional SCs, when a quasiparticle feels a different phase of the pair potential, we can expect the inner gap states localized near the surface of SCs called surface Andreev bound states (SABSs)[6, 7, 8, 4] which crucially influences the charge transport of unconventional SC junctions [4, 9]. It is known that spin-singlet $d$-wave SCs or spin-triplet $p$-wave SCs with line nodes lead to a zero-bias conductance peak (ZBCP) in normal metal/SC junctions [10, 11, 4], and it is caused by the existence of the flat bands of SABSs at zero-energy [12, 13, 14, 15]. The physical origin of ZBCP is the perfect Andreev reflection with probability unity [10, 4], and this process is equivalent to the tunneling via Majorana zero mode in topological superconductors [16, 17, 18]. ZBCP has been observed experimentally in cuprates [19, 20, 21, 22, 23, 24, 25, 26], Sr2RuO4 [27], heavy fermion CeCoIn5 [28], PuCoGa5 [29], UBe13 [30], UTe2 [31], PrOs4Sb12 [32], surface superconductivity AuSn4 [33], CaAg1-xPdx[34], CuxBi2Se3 [35], La3Ni2O7 [36], and RbCr3As3 [37]. Also, there have been studies about the charge transport of normal metal/unconventional SC junctions where the SABSs have dispersions like chiral $p$-wave [38, 39] chiral $d$-wave [40] and helical $p$-wave [41, 42]. In the case of dispersive SABS, the Andreev reflection and resulting conductance are sensitive to the Fermi surface of both normal metal and SCs [43].
@@
Recently, new type of magnets so-called “altermagnets” (AMs) was discovered [53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72]. As an extension of AMs, $p$-wave unconventional magnets (PUMs) were also proposed [73, 74]. There are many studies of AMs and PUMs exploring new phenomena in superconducting junctions with AMs and PUMs and relevant topics  [75] e.g. Andreev reflection [76, 77, 78], Josephson effect [79, 80, 81, 82, 83, 84, 85, 86, 87], topological superconductivity [88, 89, 90, 91, 92], superconducting diode effect [93, 94, 95, 96], the finite-momentum Cooper pair [97, 98, 99, 100], and so on [101, 102, 103, 104, 105, 106, 107]. In Ref. [108], the tunneling conductance in PUM/SC junctions was calculated. They found the suppression of Andreev bound states by the $p$-wave-like Fermi surface splitting in PUMs. As a result, the zero-bias conductance shows a non-monotonic change as a function of the exchange energy of the PUM order. However, in Ref. [108], they only considered the shift of the Fermi surfaces, and its Hamiltonian does not break the time-reversal symmetry, which is essentially equivalent to that of the persistent helix model  [109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 74]. Because PUMs are caused by the noncollinear spin structure, the electronic structure of PUMs does not preserve the time-reversal symmetry [73]. Thus, it is necessary to assess whether several theoretical predictions in an effective PUM model [108] are justified or not by a model taking into account the time-reversal symmetry breaking.
@@
In this paper, we investigate the tunneling conductance in superconducting junctions with PUMs. We adopt the model Hamiltonian of PUMs with time-reversal symmetry breaking suggested in Ref. [74]. Here, we focus on the junction along the $y$-direction in the low transparency limit. For spin-singlet and spin-triplet SC junctions preserving time-reversal symmetry, although the range of the transverse momentum that contributes to tunneling conductance is restricted by the spin-split feature on the Fermi surface in PUMs, the line shape of the tunneling conductance is not qualitatively affected by the PUM order. In the helical $p$-wave SC junction, the tunneling conductance is asymmetric for $eV$ with the bias voltage $V$ owing to the lack of helical edge states contributing to the conductance caused by the spin-split feature of the Fermi surface in PUM. In chiral $d$ and chiral $p$-wave SC junctions, the spin-resolved tunneling conductance has a different value for spin sectors caused by the time-reversal symmetry breaking in superconductors. Our calculations based on Ref. [74], taking account of the time-reversal symmetry breaking, qualitatively reproduce the behavior of the tunneling conductance obtained by the simplified model in Ref. [108].
@@
The construction of this paper is as follows. In Sect. II, we show the model Hamiltonian and the procedure for the calculation of the tunneling conductance in PUM/SC junctions. In Sect. III, we demonstrate the tunneling conductance and the momentum-resolved one for spin-singlet and spin-triplet pair potentials including helical $p$-wave and chiral $d$ and $p$-wave superconductivity. In Sect. IV, we conclude this work.
@@
Figure 1: (a) Image of $p$-wave unconventional magnet (PUM)/superconductor (SC) junctions. We suppose the periodic boundary condition along the $x$-direction and semi-finite systems along the $y$-direction. We consider two orbital (sublattice) degrees of freedom: orbital A (red) and B (blue) in PUMs and the single-orbital model in SCs (green). To correspond to the $p_{y}$-wave magnet/SC junction with the interface along the $x$-direction, we consider a junction along the $y$-direction. Black dotted lines indicate the tunneling between A (red) and B (blue) of PUM and SC (green). 0 and 1 indicate the indices in the Green’s function of Sect. II. The Fermi surface of PUM and that in the normal state of SC are shown in (b) and (c). (b) $\mu_{p}=-3.85t$ and $t_{J}=0.25t$, and (c) $\mu_{\mathrm{sc}}=-2.50t$. The black-dashed lines indicate the overlapping momenta $k_{x1}$ and $k_{x2}$ between PUM and SC.
@@ II Model and Methods
^$\hat{\mathcal{H}}_{\mathrm{PUM}}$ $=\sum_{\bm{k}}\hat{B}^{\dagger}\hat{H}_{\mathrm{PUM}}(\bm{k})\hat{B},$ (13)
@@
$=-\mu_{p}\hat{s}_{0}\otimes\hat{\sigma}_{0}-2t_{p}\left[\cos\frac{k_{x}}{2}\hat{s}_{0}\otimes\hat{\sigma}_{1}+\cos{k_{y}}\hat{s}_{0}\otimes\hat{\sigma}_{0}\right]$
$+2t_{J}\left[\sin\frac{k_{x}}{2}\hat{s}_{1}\otimes\hat{\sigma}_{2}+\cos{k_{y}}\hat{s}_{2}\otimes\hat{\sigma}_{3}\right],$ (14)
where $t_{p}=t$ is the nearest neighbor hopping integral in PUM, $t_{J}$ is the strength of the PUM order, and $\hat{\sigma}_{0,1,2,3}$ are the Pauli matrices in orbital (sublattice) space, respectively. This PUM model has two orbital (sublattice) degrees of freedom: orbital (sublattice) $A$ and $B$. If $t_{J}\neq 0$, the electric structure holds the spin splitting of the PUM order; otherwise, a spin-degenerate Fermi surface is obtained.
@@
In the present study, we consider the junctions along the $y$-direction shown in Fig. 1 (a) corresponding to $p_{y}$-wave UM/SC junctions along the $x$-direction. Then, we suppose the periodic boundary condition along the $x$-direction and the semi-infinite systems along the $y$-direction, where the translational invariance is broken. The Fermi surfaces on two sides of the junctions are chosen as $\mu_{\mathrm{sc}}=-2.5t$ and $\mu_{p}=-3.85t$ [Fig. 1 (b) and Fig. 1(c)]. Since the the tunneling conductance is determined by the overlapping area of Fermi surfaces between two sides of the junctions for for fixed $k_{x}$, we show how two Fermi surfaces of UPM and SC with $k_{x}$ in Fig. 1 (b) and Fig. 1 (c). Only up-spin (down-spin) configuration on the Fermi surface in PUM appears for $-k_{x2}\leq k_{x}\leq-k_{x1}$ ($k_{x1}\leq k_{x}\leq k_{x2}$) shown in Fig. 1 (b) and Fig. 1 (c). These $k_{x1}$ and $k_{x2}$ changes by $t_{J}$ shown in Fig. 2. Because $k_{x2}$ becomes larger for $t_{J}$, the Fermi surface of PUMs becomes larger with the increase of $t_{J}$.
@@
Figure 6: Tunneling conductance $\bar{\sigma}(eV)$ for spin-triplet SC junctions. (a)(d) helical $p$-wave, (b)(e) $p_{x}$-wave, and (c)(f) $p_{y}$-wave SC junctions at (a)(b)(c) $t_{J}=0$ and (d)(e)(f) $t_{J}=0.25t$. This tunneling conductance is normalized by its value in the normal state at $eV=0$: $\sigma_{\mathrm{N}}$. We choose the parameters as $U_{\mathrm{b}}=5t$, $\Delta=0.001t$, and $\delta=0.01\Delta$.
@@
in the basis $[b^{\dagger}_{A\uparrow},b^{\dagger}_{B\uparrow},b^{\dagger}_{A\downarrow},b^{\dagger}_{B\downarrow}]$ and $[c_{\uparrow},c_{\downarrow}]$. Here, $\hat{\tau}_{0,1,2,3}$ denote the Pauli matrices in Nambu space.
@@
[…]
@@ III.1 Spin-singlet superconductor junctions with time-reversal symmetry
First, we focus on the spin-singlet pair potential. In Fig. 3, we plot the momentum-resolved tunneling conductance for $\sigma(eV,k_{x})$ and spin-resolved ones $\sigma_{\uparrow}(eV,k_{x})$ and $\sigma_{\downarrow}(eV,k_{x})$ for $s$-wave, $d_{x^{2}-y^{2}}$-wave, and $d_{xy}$-wave superconductor junctions in Fig. 3 (a)(d)(g), Fig. 3 (b)(e)(h), and Fig. 3 (c)(f)(i), respectively. We see the symmetric $\sigma(eV,k_{x})$ as a function for both $eV$ and $k_{x}$ for all spin-singlet states [Figs. 3(a)(b)(c)]. The obtained $\sigma(eV,k_{x})$ are qualitatively the same as those in normal metal/SC junctions: full-gapped state of the $s$-wave [Fig. 3 (a)], the nodal structure of the $d_{x^{2}-y^{2}}$-wave [Fig. 3 (b)] and the ZBCP originating from the zero-energy SABSs of the $d_{xy}$-wave [Fig. 3 (c)]. Tunneling conductance with up-spin $\sigma_{\uparrow}(eV,k_{x})$ shown in Figs. 3 (d)(e)(f) and with down spin $\sigma_{\uparrow}(eV,k_{x})$ shown in Figs. 3 (g)(h)(i) are not symmetric for $k_{x}$ and we find the relation: $\sigma_{\uparrow}(eV,k_{x})=\sigma_{\downarrow}(eV,-k_{x})$. Then the range of $k_{x}$ contributing to the conductance process with up-spin is restricted for $-k_{x2}\leq k_{x}\leq k_{x1}$ with $\sigma_{\uparrow}(eV,k_{x})$ [Figs. 3 (d)(e)(f)], while, with down spin for $-k_{x1}\leq k_{x}\leq k_{x2}$ with $\sigma_{\downarrow}(eV,k_{x})$ [Figs. 3 (g)(h)(i)]. This originates from only one spin configuration of the Fermi surface in PUMs available for charge transport shown in Fig. 1 (b): for $-k_{x2}\leq k_{x}\leq-k_{x1}$ with up-spin and $k_{x1}\leq k_{x}\leq k_{x2}$ with down-spin.

### v2

# Tunneling conductance in superconducting junctions with $p$-wave unconventional magnets breaking time-reversal symmetry

@@ Abstract
A new type of magnet called $p$-wave unconventional magnet is proposed, stimulated by the discovery of altermagnet. We study the tunneling conductance of $p$-wave unconventional magnet/superconductor junctions by adopting the effective Hamiltonian of $p$-wave unconventional magnets with time-reversal symmetry breaking, suggested in Ref [arXiv: 2309.01607 (2024)]. The tunneling conductance shows an asymmetric behavior with respect to bias voltage in the helical $p$-wave superconductor junctions. It is caused by the missing of helical edge states contributing to the charge conductance owing to the momentum-dependent spin-split feature of the Fermi surface in $p$-wave unconventional magnets. In chiral $d$ and $p$-wave superconductor junctions, the resulting spin-resolved tunneling conductance takes a different value for spin sectors due to the time-reversal symmetry breaking in superconductors. Our results qualitatively reproduce the results based on the simplified Hamiltonian in Ref [J. Phys. Soc. Jpn. 93, 114703 (2024)], where only the odd function of the exchange coupling of $p$-wave unconventional magnets is taken into account, which gives the shift of the Fermi surface and preserves the time-reversal symmetry similar to the spin-orbit coupling.
@@ I Introduction
Charge transport of superconducting junctions is a fundamental problem of superconductivity and mesoscopic physics due to the presence of Andreev reflection [1, 2]. Andreev reflection is a unique process in inhomogeneous superconductors (SCs), like superconducting junctions, where an injected electron (hole) is reflected as a hole (electron) at the interface, and it can generate the inner gap states called Andreev bound states [3, 4, 5]. In unconventional SCs, when a quasiparticle feels a different phase of the pair potential, we can expect the inner gap states localized near the surface of SCs called surface Andreev bound states (SABSs)[6, 7, 8, 4] which crucially influences the charge transport of unconventional SC junctions [4, 9]. It is known that spin-singlet $d$-wave SCs or spin-triplet $p$-wave SCs with line nodes lead to a zero-bias conductance peak (ZBCP) in normal metal/SC junctions [10, 11, 4], and it is caused by the existence of the flat bands of SABSs at zero-energy [12, 13, 14, 15]. The physical origin of ZBCP is the perfect Andreev reflection with probability unity [10, 4], and this process is equivalent to the tunneling via Majorana zero mode in topological superconductors [16, 17, 18]. ZBCP has been observed experimentally in cuprates [19, 20, 21, 22, 23, 24, 25, 26], Sr2RuO4 [27], heavy fermion CeCoIn5 [28], PuCoGa5 [29], UBe13 [30], UTe2 [31], PrOs4Sb12 [32], surface superconductivity AuSn4 [33], CaAg1-xPdxP [34], CuxBi2Se3 [35], La3Ni2O7 [36], and RbCr3As3 [37]. Also, there have been studies about the charge transport of normal metal/unconventional SC junctions where the SABSs have dispersions like chiral $p$-wave [38, 39] chiral $d$-wave [40] and helical $p$-wave [41, 42]. In the case of dispersive SABS, the Andreev reflection and resulting conductance are sensitive to the Fermi surface of both the normal metal and SCs [43].
@@
Recently, new type of magnets so-called “altermagnets” (AMs) was discovered [53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73]. As an extension of AMs, $p$-wave unconventional magnets (PUMs) were also proposed [74, 75, 76] and observed experimentally in NiI2 [77]. There are many studies of AMs and PUMs exploring new phenomena in superconducting junctions with AMs and PUMs and relevant topics [73] e.g. Andreev reflection [78, 79, 80], Josephson effect [81, 82, 83, 84, 85, 86, 87, 88, 89], topological superconductivity [90, 91, 92, 93, 94], superconducting diode effect [95, 96, 97, 98], the finite-momentum Cooper pair [99, 100, 101, 102], and so on [103, 104, 105, 106, 107, 108, 109]. In Ref. [110], the tunneling conductance in PUM/SC junctions was calculated by using the simplified continuum model. They found the suppression of Andreev bound states by the $p$-wave-like Fermi surface splitting in PUMs. As a result, the zero-bias conductance shows a non-monotonic change as a function of the exchange energy of the PUM order. However, in Ref. [110], they only considered the shift of the Fermi surfaces in the continuum model, and its Hamiltonian does not break the time-reversal symmetry, which is essentially equivalent to that of the persistent helix model [111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 75]. Because PUMs are caused by the noncollinear spin structure, the electronic structure of PUMs does not preserve the time-reversal symmetry [74, 75]. While, the energy spectrum in the momentum space satisfies the energy dispersion: $E(\bm{k},\sigma)=E(-\bm{k},-\sigma)$ with the momentum $\bm{k}$ and the spin $\sigma$, owing to $T\tau$-symmetry in Ref. [74] and $C_{2}||t]-$symmetry in Ref. [75]. In junctions with PUM, the spin structure near the surface/interface might affect transport properties. By adopting the tight-binding model compared with the continuum model without the microscopic structure in the real space [110], we can deal with not only the effective model of PUM [75] but also the spin structure of PUM at the interface. Thus, it is necessary to assess whether several theoretical predictions in an effective PUM model [110] are justified or not by a model taking into account the time-reversal symmetry breaking.
@@
In this paper, we investigate the tunneling conductance in superconducting junctions with PUMs. We adopt the model Hamiltonian of PUMs with time-reversal symmetry breaking suggested in Ref. [75]. For using this Hamiltonian, we adopt the tight-binding model to focus on the spin structure of PUMs at the interface, and we choose the junction along the $y$-direction in the low transparency limit. For spin-singlet and spin-triplet SC junctions preserving time-reversal symmetry, although the range of the transverse momentum that contributes to tunneling conductance is restricted by the spin-split feature on the Fermi surface in PUMs, the line shape of the tunneling conductance is not qualitatively affected by the PUM order. In the helical $p$-wave SC junction, the tunneling conductance is asymmetric for $eV$ with the bias voltage $V$ owing to the missing of helical edge states contributing to the conductance caused by the spin-split feature of the Fermi surface in PUM. In chiral $d$ and chiral $p$-wave SC junctions, the spin-resolved tunneling conductance has a different value for spin sectors caused by the time-reversal symmetry breaking in superconductors. Our calculations based on Ref. [75], taking account of the time-reversal symmetry breaking, qualitatively reproduce the behavior of the tunneling conductance obtained by the simplified model in Ref. [110].
@@
The construction of this paper is as follows. In Sect. II, we show the model Hamiltonian and the procedure for the calculation of the tunneling conductance in PUM/SC junctions. In Sect. III, we demonstrate the tunneling conductance and the momentum-resolved one for spin-singlet and spin-triplet pair potentials, including helical $p$-wave and chiral $d$ and $p$-wave superconductivity. In Sect. IV, we conclude this work.
@@
Figure 1: (a) Image of $p$-wave unconventional magnet (PUM)/superconductor (SC) junctions. We suppose the periodic boundary condition along the $x$-direction and semi-infinite systems along the $y$-direction. We consider two orbital (sublattice) degrees of freedom: orbital A (red) and B (blue) in PUMs and the single-orbital model in SCs (green). To correspond to the $p_{y}$-wave magnet/SC junction with the interface along the $x$-direction, we consider a junction along the $y$-direction. Black dotted lines indicate the tunneling between A (red) and B (blue) of PUM and SC (green). 0 and 1 indicate the indices in the Green’s function of Sect. II. The Fermi surface of PUM and that in the normal state of SC are shown in (b) and (c). (b) $\mu_{p}=-3.85t$ and $t_{J}=0.25t$, and (c) $\mu_{\mathrm{sc}}=-2.50t$. The black-dashed lines indicate the overlapping momenta $k_{x1}$ and $k_{x2}$ between PUM and SC.
@@ II Model and Methods
$\hat{\mathcal{H}}_{\mathrm{PUM}}$ $=\sum_{\bm{k}}\hat{B}^{\dagger}\hat{H}_{\mathrm{PUM}}(\bm{k})\hat{B}$, (13)
@@
[…]
where $t_{p}=t$ is the nearest neighbor hopping integral in PUM, $t_{J}$ is the strength of the PUM order, and $\hat{\sigma}_{0,1,2,3}$ are the Pauli matrices in orbital (sublattice) space, respectively. This PUM model has two orbital (sublattice) degrees of freedom: orbital (sublattice) $A$ and $B$. For $t_{J}\neq 0$, $C_{2}||t]-$symmetry is preserved, even though time-reversal symmetry is broken. If $t_{J}\neq 0$, the electric structure holds the spin splitting of the PUM order; otherwise, a spin-degenerate Fermi surface is obtained.
@@
In the present study, we consider the junctions along the $y$-direction shown in Fig. 1 (a) corresponding to $p_{y}$-wave UM/SC junctions along the $x$-direction. Then, we suppose the periodic boundary condition along the $x$-direction and the semi-infinite systems along the $y$-direction, where the translational invariance is broken. The Fermi surfaces on two sides of the junctions are chosen as $\mu_{\mathrm{sc}}=-2.5t$ and $\mu_{p}=-3.85t$ [Fig. 1 (b) and Fig. 1(c)]. Since the tunneling conductance is determined by the overlapping area of Fermi surfaces between two sides of the junctions for for fixed $k_{x}$, we show how two Fermi surfaces of UPM and SC with $k_{x}$ in Fig. 1 (b) and Fig. 1 (c). Only up-spin (down-spin) configuration on the Fermi surface in PUM appears for $k_{x2}\leq k_{x}\leq-k_{x1}$ ($k_{x1}\leq k_{x}\leq k_{x2}$) shown in Fig. 1 (b) and Fig. 1 (c). These $k_{x1}$ and $k_{x2}$ changes by $t_{J}$ shown in Fig. 2. Because $k_{x2}$ becomes larger for $t_{J}$, the Fermi surface of PUMs becomes larger with the increase of $t_{J}$.
@@
Figure 6: Tunneling conductance $\bar{\sigma}(eV)$ for spin-triplet SC junctions. (a)(d) helical $p$-wave, (b)(e) $p_{x}$-wave, and (c)(f) $p_{y}$-wave SC junctions at (a)(b)(c) $t_{J}=0$ and (d)(e)(f) $t_{J}=0.25t$. In (d), we also plot $\bar{\sigma}(eV)-\bar{\sigma}(eV)]/\sigma_{\mathrm{N}}$ to show the asymmetry of the conductance for $eV$. The tunneling conductance is normalized by its value in the normal state at $eV=0$: $\sigma_{\mathrm{N}}$. We choose the parameters as $U_{\mathrm{b}}=5t$, $\Delta=0.001t$, and $\delta=0.01\Delta$.
@@
in the basis $b^{\dagger}_{A\uparrow},b^{\dagger}_{B\uparrow},b^{\dagger}_{A\downarrow},b^{\dagger}_{B\downarrow}]$ and $c_{\uparrow},c_{\downarrow}]$. Here, $\hat{\tau}_{0,1,2,3}$ denote the Pauli matrices in Nambu space.
@@
where $\mathrm{Tr^{\prime}_{\uparrow\uparrow}}$ and $\mathrm{Tr^{\prime}_{\downarrow\downarrow}}$ are the trace onto only $\uparrow\uparrow$ and $\downarrow\downarrow$ in the electron space, respectively. In the next section, for each pair potential, we calculate $\bar{\sigma}(eV)$, $\bar{\sigma}_{\uparrow}(eV)$, and $\bar{\sigma}_{\downarrow}(eV)$ for each momentum $k_{x}$ and the bias $eV$, and we also show the tunneling conductance. Here, we normalized the tunneling conductance by using $\sigma_{\mathrm{N}}$ with the tunneling conductance in the normal state at zero bias voltage $eV=0$.
@@ III.1 Spin-singlet superconductor junctions with time-reversal symmetry
First, we focus on the spin-singlet pair potential. In Fig. 3, we plot the momentum-resolved tunneling conductance for $\sigma(eV,k_{x})$ and spin-resolved ones $\sigma_{\uparrow}(eV,k_{x})$ and $\sigma_{\downarrow}(eV,k_{x})$ for $s$-wave, $d_{x^{2}-y^{2}}$-wave, and $d_{xy}$-wave superconductor junctions in Fig. 3 (a)(d)(g), Fig. 3 (b)(e)(h), and Fig. 3 (c)(f)(i), respectively. We see the symmetric $\sigma(eV,k_{x})$ as a function for both $eV$ and $k_{x}$ for all spin-singlet states [Figs. 3(a)(b)(c)]. The obtained $\sigma(eV,k_{x})$ are qualitatively the same as those in normal metal/SC junctions: full-gapped state of the $s$-wave [Fig. 3 (a)], the nodal structure of the $d_{x^{2}-y^{2}}$-wave [Fig. 3 (b)] and the ZBCP originating from the zero-energy SABSs of the $d_{xy}$-wave [Fig. 3 (c)]. Tunneling conductance with up-spin $\sigma_{\uparrow}(eV,k_{x})$ shown in Figs. 3 (d)(e)(f) and with down spin $\sigma_{\downarrow}(eV,k_{x})$ shown in Figs. 3 (g)(h)(i) are not symmetric for $k_{x}$ and we find the relation: $\sigma_{\uparrow}(eV,k_{x})=\sigma_{\downarrow}(eV,-k_{x})$. Then the range of $k_{x}$ contributing to the conductance process with up-spin is restricted for $k_{x2}\leq k_{x}\leq k_{x1}$ with $\sigma_{\uparrow}(eV,k_{x})$ [Figs. 3 (d)(e)(f)], while, with down spin for $k_{x1}\leq k_{x}\leq k_{x2}$ with $\sigma_{\downarrow}(eV,k_{x})$ [Figs. 3 (g)(h)(i)]. This originates from only one spin configuration of the Fermi surface in PUMs available for charge transport shown in Fig. 1 (b): for $k_{x2}\leq k_{x}\leq-k_{x1}$ with up-spin and $k_{x1}\leq k_{x}\leq k_{x2}$ with down-spin.